# Resonances and resonance expansions for $\delta$-interactions on the half-space

### v2

@@ 6. Resonance Expansion for the Wave and Schrödinger Equation with Point Interaction
By the functional calculus, the propagator of Eq. 6.1 can be written as
(6.8) $U(t)=\int_{0}^{+\infty}\frac{\sin{tz}}{z}\mathop{}\!\mathrm{d}E_{z}+\frac{\sin{t\lambda}}{\lambda}\biggl{(}\cdot,G_{\sqrt{\lambda},y}^{\textup{D}}\biggr{)}G_{\sqrt{\lambda},y}^{\textup{D}}.$
If $\lambda=0$ is an eigenvalue (case $ii)$ in the statement), the last term is replaced by
@@
Since for $z$ near $\lambda\neq 0$, $R(z)=(\lambda^{2}-z^{2})^{-1}\biggl{(}\cdot,G_{\sqrt{\lambda},y}^{\textup{D}}\biggr{)}G_{\sqrt{\lambda},y}^{\textup{D}}+Q(z)$, where $Q(z)$ is holomorphic near $\lambda$ and $\operatorname{Res}_{z=\lambda}{(\lambda^{2}-z^{2})}=-(2\lambda)^{-1}$, we have
(6.9) $\frac{\sin{t\lambda}}{\lambda}\biggl{(}\cdot,G_{\sqrt{\lambda},y}^{\textup{D}}\biggr{)}G_{\sqrt{\lambda},y}^{\textup{D}}=-\operatorname{Res}_{z=\lambda}{(iR(z)e^{-izt})}-\operatorname{Res}_{z=-\lambda}{(iR(-z)e^{-izt})}=\Pi_{\lambda}+\Pi_{-\lambda}.$
If $\lambda=0$, instead
(6.10) $t\,\biggl{(}\cdot,G_{0,y}^{\textup{D}}\biggr{)}G_{0,y}^{\textup{D}}=-t\operatorname{Res}_{z=0}{(iR(z))}.$
In the following, the term for $\lambda=0$ will not be reported whenever it is not needed to distinguish the two cases.
@@
(6.11) $\rho\Bigl{(}R(z)-R(-z)\Bigr{)}\rho w_{1}=\rho\Bigl{(}R(z)-R(-z)\Bigr{)}(1+z^{2})^{-1}\biggl{(}1-\Delta_{\alpha,y}^{\textup{D}}\biggr{)}\rho w_{1}.$
@@
The resolvent is holomorphic in a neighborhhod of the origin and hence the integral over $\sigma_{\varepsilon}$ converges to 00 as $\varepsilon\to 0^{+}$, unless zero is an eigenvalue for $-\Delta_{\alpha,y}^{\textup{D}}$. In that case, we study the expression $R(z)-R(-z)$ near zero. We recall that 00 is an eigenvalue if and only if $\alpha=-\frac{1}{8\pi y_{3}}$. In this case
@@
$\omega_{r}=\Set{s-i\,\Bigl{(}A+\varepsilon+\delta\ln{(1+\left\lvert s\right\rvert)}\Bigr{)}}{s\in[-r,r]}$
$\gamma_{r}^{\pm}=\Set{\pm r-is}{0\leq s\leq A+\varepsilon+\delta\ln{(1+r)}},\quad\gamma_{r}=\gamma_{r}^{+}\cup\gamma_{r}^{-}$
@@
$\Omega_{A}=\Set{z}{\operatorname{Im}{z}\geq-A-\varepsilon-\delta\ln{(1+\left\lvert\operatorname{Re}{z}\right\rvert)}}\setminus\Set{0}$
@@
$\Pi^{\prime}_{A}(t)=\sum_{\Set{z_{j}\in\Omega_{A}}\cup\Set{z_{j}=\pm\lambda}}\operatorname{Res}_{z=z_{j}}{\Bigl{(}\rho R(z)\rho e^{-izt}\Bigr{)}}.$
By combining Eq. 6.8, Eq. 6.9 (or Eq. 6.10 when zero is an eigenvalue) and the residue theorem, we can write that
(6.12) $\rho U(t)\rho=\Pi^{\prime}_{A}(t)+E_{\omega_{r}}(t)+E_{\gamma_{r}}(t)+E_{\gamma_{r}^{\infty}}(t)+\Pi_{\lambda}+\Pi_{-\lambda},$
@@
We note that $\Pi_{\lambda}+\Pi_{-\lambda}$ cancels out with the term in $\Pi^{\prime}_{A}$ containing the residue at $z=\pm\lambda$. So, by putting together this cancellations in the new term, $\Pi_{A}$, Eq. 6.12 becomes
@@
Now we prove that the the last two terms in Eq. 6.12 are negligible in the limit $r\to+\infty$. To do so, we consider $w_{1}\in H^{2}(K)$. This allows us to use Eq. 6.11 to Eq. 6.3 with $j=1$ and obtain the following bound
@@
($\omega$ is the curve $\omega=\Set{s-i\,\Bigl{(}A+\varepsilon+\delta\ln{(1+\left\lvert s\right\rvert)}\Bigr{)}}{s\in\mathbb{R}}$).
@@
(6.13) $\left\lVert E_{\omega}(t)w_{1}\right\rVert_{H^{2}(K)}\leq Ce^{-tA}\left\lVert w_{1}\right\rVert_{L^{2}(K)}.$
In order to do so, we use Eq. 6.3 with $j=2$ and the assumption that there exist a compact set containing $\omega$ in which there are no poles of $R(z)$. We obtain
@@
This last integral is finite if $-(t-T)\delta+1<-1$, which is equivalent to $t>T+\frac{2}{\delta}$.
In Eq. 6.13 the above bound is in terms of the $L^{2}(K)$ norm of $w_{1}$ for $w_{1}\in H^{2}(K)$. Since, $H^{2}(K)$ is dense in $L^{2}(K)$, then the same bound holds for any $w_{1}\in L^{2}(K)$ and the theorem holds for any $w_{1}\in L^{2}(K)$.
@@
(6.14) $\begin{split}e^{-itH_{+}}(x,x^{\prime})&=(2\pi it)^{-\frac{3}{2}}e^{\frac{i}{2t}\left\lvert x-x^{\prime}\right\rvert^{2}}+\sum_{n}2z_{n}e^{-itz_{n}^{2}}\operatorname{Res}_{z=z_{n}}{\biggl{(}\Bigl{(}\Gamma^{\textup{D}}_{\alpha,y}(z)\Bigr{)}^{-1}\biggr{)}}G^{\textup{D}}_{z_{n},y}(x)G^{\textup{D}}_{z_{n},y}(x^{\prime})\\ &-\frac{1}{2\pi}\int_{0}^{+\infty}\biggl{(}\hat{R}\Bigl{(}e^{-i\frac{\pi}{4}}s\Bigr{)}-\hat{R}\Bigl{(}-e^{-i\frac{\pi}{4}}s\Bigr{)}\biggr{)}e^{-ts}\mathop{}\!\mathrm{d}s,\end{split}$
with $\hat{R}(z)=(H-z^{2})^{-1}-(-\Delta^{D}-z^{2})^{-1}$. The sum in Eq. 6.14 is taken over all $z_{n}$ lying in the region $-\frac{\pi}{4}<\operatorname{Arg}{k}<0$ such that $z_{n}^{2}$ is a resonance.
@@
$-\frac{1}{2\pi i}\int_{0}^{+\infty}e^{-itz^{2}}\Bigl{(}\hat{R}(z)-\hat{R}(-z)\Bigr{)}(x,x^{\prime})2z\mathop{}\!\mathrm{d}z.$
@@
(6.15) $\begin{split}\frac{1}{2\pi i}\int_{C_{R}}e^{-itz^{2}}\Bigl{(}\hat{R}(z)-\hat{R}(-z)\Bigr{)}(x,x^{\prime})2z\mathop{}\!\mathrm{d}z\ +\ &\frac{1}{2\pi i}\int_{\gamma_{R}}e^{-itz^{2}}\Bigl{(}\hat{R}(z)-\hat{R}(-z)\Bigr{)}(x,x^{\prime})2z\mathop{}\!\mathrm{d}z\\ +\ &\sum_{n}2z_{n}e^{-itz_{n}^{2}}\operatorname{Res}_{z=z_{n}}{\Bigl{(}\hat{R}(z)-\hat{R}(-z)\Bigr{)}}(x,x^{\prime}),\end{split}$
where $C_{R}$ is the arc of circumference of radius $R$ going clockwise from $z=R$ to $z=Re^{-i\pi/4}$, $\gamma_{R}$ is the segment going from $z=Re^{-i\pi/4}$ towards the origin and the sum is over the $z_{n}$ in the region delimited by $C_{R}$, $\gamma_{R}$ and the real positive semiaxis.
We prove that the integral over $C_{R}$ vanishes for $R\to+\infty$. We prove this just for the term with $\hat{R}(z)$. Analogous observations prove the same for $\hat{R}(-z)$. By Proposition 3.1, there are infinite resonances $z_{n}$ and $\left\lvert z_{n}\right\rvert\to+\infty$ for $n\to+\infty$. This means that, in order to study the limit $R\to+\infty$, we have to distinguish the cases
  • $R\neq\left\lvert z_{n}\right\rvert$, $\forall n\in\mathbb{N}$;
@@
(6.16) $\int_{C_{R}}e^{-itz^{2}}\biggl{(}\alpha-\frac{iz}{4\pi}+\frac{e^{2izy_{3}}}{8\pi y_{3}}\biggr{)}^{-1}\frac{e^{iz\left\lvert x-y\right\rvert}}{4\pi\left\lvert x-y\right\rvert}\frac{e^{iz\left\lvert x^{\prime}-y\right\rvert}}{4\pi\left\lvert x^{\prime}-y\right\rvert}2z\mathop{}\!\mathrm{d}z.$
@@
Since we are in the complex lower half-plane, $\sin{\theta}<0$. So if the other factor in the exponent is positive, then in the limit the integral vanishes. By recalling that $-\frac{\pi}{4}<\theta<0$, we have that
@@
which is positive for $t>\frac{\sqrt{2}}{2R}\Bigl{(}\left\lvert x-y\right\rvert+\left\lvert x^{\prime}-y\right\rvert\Bigr{)}$ and hence the integral vanishes in the limit.
In the other case, the path of integration has to be modified because the previous path would pass across the pole $z_{n}$. So, the integral Eq. 6.16 has to be replaced by the following sum of integrals of the same function
$-\frac{1}{2\pi i}\int_{C_{R}^{1}}\dots\mathop{}\!\mathrm{d}z-\frac{1}{2\pi i}\int_{C_{\varepsilon}}\dots\mathop{}\!\mathrm{d}z-\frac{1}{2\pi i}\int_{C_{R}^{2}}\dots\mathop{}\!\mathrm{d}z,$
@@
$-\frac{1}{2\pi i}\int_{C_{\varepsilon}}\dots\mathop{}\!\mathrm{d}z=-\frac{1}{2}\operatorname{Res}_{z=z_{n}}{\Biggl{(}e^{-itz^{2}}\biggl{(}\alpha-\frac{iz}{4\pi}+\frac{e^{2izy_{3}}}{8\pi y_{3}}\biggr{)}^{-1}\frac{e^{iz\left\lvert x-y\right\rvert}}{4\pi\left\lvert x-y\right\rvert}\frac{e^{iz\left\lvert x^{\prime}-y\right\rvert}}{4\pi\left\lvert x^{\prime}-y\right\rvert}2z\Biggr{)}}.$
@@
$\begin{split}\left\lvert-\frac{1}{2\pi i}\int_{C_{\varepsilon}}\dots\mathop{}\!\mathrm{d}z\right\rvert&=\left\lvert e^{-itz_{n}^{2}}\frac{4\pi i}{e^{2iy_{3}z_{n}}-1}\frac{e^{iz_{n}\left\lvert x-y\right\rvert}}{4\pi\left\lvert x-y\right\rvert}\frac{e^{iz_{n}\left\lvert x^{\prime}-y\right\rvert}}{4\pi\left\lvert x^{\prime}-y\right\rvert}z_{n}\right\rvert\\ &\leq C^{\prime}(x,x^{\prime},y)e^{r_{n}\sin{\theta_{n}}(2tr_{n}\cos{\theta_{n}}-\left\lvert x-y\right\rvert-\left\lvert x^{\prime}-y\right\rvert)}.\end{split}$
@@
with the last expression being positive for $t>\frac{\sqrt{2}}{2r_{n}}\Bigl{(}\left\lvert x-y\right\rvert+\left\lvert x^{\prime}-y\right\rvert\Bigr{)}$. So, in the limit $n\to+\infty$ also the integral on $C_{\varepsilon}$ vanishes for $t$ large enough.
Going back to Eq. 6.15, we consider the sum in it. First we observe that $\hat{R}(-z)$ has no poles in the region considered, because all its poles are in the upper complex plane. The expression of $(-\Delta_{\alpha,y}^{\textup{D}}-z^{2})^{-1}$ and the fact that for $R\to+\infty$ the region of interest becomes $-\frac{\pi}{4}<\arg{z}<0$ let us recollect the expression present in Eq. 6.14.

### v3

@@ 6. Resonance Expansion for the wave and Schrödinger Equation with Point Interaction
By the functional calculus, the propagator of Eq.˜6.1 can be written as
$U(t)=\int_{0}^{+\infty}\frac{\sin{tz}}{z}\mathop{}\!\mathrm{d}E_{z}+\frac{\sin{t\mu}}{\mu}\biggl(\cdot,G_{\mu,y}^{\textup{D}}\biggr)G_{\mu,y}^{\textup{D}},$ (6.8)
where $\mu^{2}$ is the possible eigenvalue of $\Delta_{\alpha,y}^{\textup{D}}$ (the term is not present when there is no eigenvalue).If the eigenvalue is zero (case ii)ii) in the statement), the last term is replaced by
@@
Since for $z$ near $\mu\neq 0$, $R(z)=(-\mu^{2}-z^{2})^{-1}\biggl(\cdot,G_{\mu,y}^{\textup{D}}\biggr)G_{\mu,y}^{\textup{D}}+Q(z)$, where $Q(z)$ is holomorphic near $\mu^{2}$ and $\operatorname{Res}_{z=i\mu}{(-\mu^{2}-z^{2})}=-(2\mu)^{-1}$, we have
$\frac{\sin{t\mu}}{\mu}\biggl(\cdot,G_{\mu,y}^{\textup{D}}\biggr)G_{\mu,y}^{\textup{D}}=-\operatorname{Res}_{z=i\mu}{(iR(z)e^{-izt})}-\operatorname{Res}_{z=-i\mu}{(iR(-z)e^{-izt})}=\Pi_{\mu}+\Pi_{-\mu}.$ (6.9)
If $\mu=0$, instead
$t\,\biggl(\cdot,G_{0,y}^{\textup{D}}\biggr)G_{0,y}^{\textup{D}}=-t\operatorname{Res}_{z=0}{(iR(z))}.$ (6.10)
In the following, the term for $\mu=0$ will not be reported whenever it is not needed to distinguish the two cases.
@@
$\rho\Bigl(R(z)-R(-z)\Bigr)\rho w_{1}=\rho\Bigl(R(z)-R(-z)\Bigr)(1+z^{2})^{-1}\biggl(1-\Delta_{\alpha,y}^{\textup{D}}\biggr)\rho w_{1}.$ (6.11)
@@
The resolvent is holomorphic in a neighborhhod of the origin and hence the integral over $\sigma_{\varepsilon}$ converges to 0 as $\varepsilon\to 0^{+}$, unless zero is an eigenvalue for $\Delta_{\alpha,y}^{\textup{D}}$. In that case, we study the expression R(z)$R(z)$-R(-z) near zero. We recall that 0 is an eigenvalue if and only if $\alpha=-\frac{1}{8\pi y_{3}}$. In this case
@@
[…]
@@
[…]
@@
$\Pi^{\prime}_{A}(t)=\sum_{\Set{z_{j}\in\Omega_{A}}}\operatorname{Res}_{z=z_{j}}{\Bigl(\rho R(z)\rho e^{-izt}\Bigr)}.$
By combining Eq.˜6.8, Eq.˜6.9 (or Eq.˜6.10 when zero is an eigenvalue) and the residue theorem, we can write that
$\rho U(t)\rho=\Pi^{\prime}_{A}(t)+E_{\omega_{r}}(t)+E_{\gamma_{r}}(t)+E_{\gamma_{r}^{\infty}}(t)+\Pi_{\mu}+\Pi_{-\mu},$ (6.12)
@@
We note that $\Pi_{-\mu}$ cancels out with the term in $\Pi^{\prime}_{A}$ containing the residue at $z=-\mu$. So, by putting together this cancellations in the new term, $\Pi_{A}$, Eq.˜6.12 becomes
@@
Now we prove that the the last two terms in Eq.˜6.12 are negligible in the limit $r\to+\infty$. To do so, we consider $w_{1}\in H^{2}(K)$. This allows us to use Eq.˜6.11 to Eq.˜6.3 with $j=1$ and obtain the following bound
@@
[…]
@@
$\left\lVert E_{\omega}(t)w_{1}\right\rVert_{H^{2}(K)}\leq Ce^{-tA}\left\lVert w_{1}\right\rVert_{L^{2}(K)}.$ (6.13)
In order to do so, we use Eq.˜6.3 with $j=2$ and the assumption that there exist a compact set containing $\omega$ in which there are no poles of $R(z)$. We obtain
@@
This last integral is finite if $(t-T)\delta+1<-1$, which is equivalent to $t>T+\frac{2}{\delta}$.
In Eq.˜6.13 the above bound is in terms of the $L^{2}(K)$ norm of $w_{1}$ for $w_{1}\in H^{2}(K)$. Since, $H^{2}(K)$ is dense in $L^{2}(K)$, then the same bound holds for any $w_{1}\in L^{2}(K)$ and the theorem holds for any $w_{1}\in L^{2}(K)$.
@@
$\begin{split}e^{-itH_{+}}(x,x^{\prime})&=(2\pi it)^{-\frac{3}{2}}e^{\frac{i}{2t}\left\lvert x-x^{\prime}\right\rvert^{2}}+\sum_{n}2z_{n}e^{-itz_{n}^{2}}\operatorname{Res}_{z=z_{n}}{\biggl(\Bigl(\Gamma^{\textup{D}}_{\alpha,y}(z)\Bigr)^{-1}\biggr)}G^{\textup{D}}_{z_{n},y}(x)G^{\textup{D}}_{z_{n},y}(x^{\prime})\\ &-\frac{1}{2\pi}\int_{0}^{+\infty}\biggl(\hat{R}\Bigl(e^{-i\frac{\pi}{4}}s\Bigr)-\hat{R}\Bigl(-e^{-i\frac{\pi}{4}}s\Bigr)\biggr)e^{-ts}\mathop{}\!\mathrm{d}s,\end{split}$ (6.14)
with $\hat{R}(z)=(H-z^{2})^{-1}-(-\Delta^{D}-z^{2})^{-1}$. The sum in Eq.˜6.14 is taken over all resonances $z_{n}$ with $\frac{\pi}{4}<\operatorname{Arg}{z_{n}}<0$.
@@
$\frac{1}{2\pi i}\int_{0}^{+\infty}e^{-itz^{2}}\Bigl(\hat{R}(z)-\hat{R}(-z)\Bigr)(x,x^{\prime})2z\mathop{}\!\mathrm{d}z.$
@@
$\begin{split}\frac{1}{2\pi i}\int_{C_{R}}e^{-itz^{2}}\Bigl(\hat{R}(z)-\hat{R}(-z)\Bigr)(x,x^{\prime})2z\mathop{}\!\mathrm{d}z\ +\ &\frac{1}{2\pi i}\int_{\gamma_{R}}e^{-itz^{2}}\Bigl(\hat{R}(z)-\hat{R}(-z)\Bigr)(x,x^{\prime})2z\mathop{}\!\mathrm{d}z\\ +\ &\sum_{n}2z_{n}e^{-itz_{n}^{2}}\operatorname{Res}_{z=z_{n}}{\Bigl(\hat{R}(z)-\hat{R}(-z)\Bigr)}(x,x^{\prime}),\end{split}$ (6.15)
where $C_{R}$ is the arc of circumference of radius $R$ going clockwise from $z=R$ to $z=Re^{-i\pi/4}$, $\gamma_{R}$ is the segment going from $z=Re^{-i\pi/4}$ towards the origin and the sum is over the $z_{n}$ in the region delimited by $C_{R}$, $\gamma_{R}$ and the real positive semi-axis.
We prove that the integral over $C_{R}$ vanishes for $R\to+\infty$. We prove this just for the term with $\hat{R}(z)$. Analogous observations prove the same for $\hat{R}(-z)$. By Proposition˜3.4, there are infinite resonances $z_{n}$ and $\left\lvert z_{n}\right\rvert\to$$\infty$ for $n\to+\infty$. This means that, in order to study the limit $R\to+\infty$, we have to distinguish the cases
  • $R\neq\left\lvert z_{n}\right\rvert$, for all $n\in\mathbb{N}$;
@@
$\int_{C_{R}}e^{-itz^{2}}\biggl(\alpha-\frac{iz}{4\pi}+\frac{e^{2izy_{3}}}{8\pi y_{3}}\biggr)^{-1}\frac{e^{iz\left\lvert x-y\right\rvert}}{4\pi\left\lvert x-y\right\rvert}\frac{e^{iz\left\lvert x^{\prime}-y\right\rvert}}{4\pi\left\lvert x^{\prime}-y\right\rvert}2z\mathop{}\!\mathrm{d}z.$ (6.16)
@@
Since we are in the complex lower half-plane, $\sin{\theta}<0$. So if the other factor in the exponent is positive, then in the limit the integral vanishes. By recalling that $\frac{\pi}{4}<\theta<0$, we have that
@@
[…]
In the other case, the path of integration has to be modified because the previous path would pass across the pole $z_{n}$. So, the integral Eq.˜6.16 has to be replaced by the following sum of integrals of the same function
$\frac{1}{2\pi i}\int_{C_{R}^{1}}\dots\mathop{}\!\mathrm{d}z-\frac{1}{2\pi i}\int_{C_{\varepsilon}}\dots\mathop{}\!\mathrm{d}z-\frac{1}{2\pi i}\int_{C_{R}^{2}}\dots\mathop{}\!\mathrm{d}z,$
@@
$\frac{1}{2\pi i}\int_{C_{\varepsilon}}\dots\mathop{}\!\mathrm{d}z=-\frac{1}{2}\operatorname{Res}_{z=z_{n}}{\Biggl(e^{-itz^{2}}\biggl(\alpha-\frac{iz}{4\pi}+\frac{e^{2izy_{3}}}{8\pi y_{3}}\biggr)^{-1}\frac{e^{iz\left\lvert x-y\right\rvert}}{4\pi\left\lvert x-y\right\rvert}\frac{e^{iz\left\lvert x^{\prime}-y\right\rvert}}{4\pi\left\lvert x^{\prime}-y\right\rvert}2z\Biggr)}.$
@@
[…]
@@
[…]
Going back to Eq.˜6.15, we consider the sum in it. First we observe that $\hat{R}(-z)$ has no poles in the region considered, because all its poles are in the upper complex plane. The expression of $(-\Delta_{\alpha,y}^{\textup{D}}-z^{2})^{-1}$ and the fact that for $R\to+\infty$ the region of interest becomes $\frac{\pi}{4}<\operatorname{Arg}{z}<0$ let us recollect the expression present in Eq.˜6.14.